# MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting

### v3

@@ Abstract
Long time series forecasting aims to utilize historical information to forecast future states over extended horizons. Traditional RNN-based series forecasting methods struggle to effectively address long-term dependencies and gradient issues in long time series problems. Recently, SegRNN has emerged as a leading RNN-based model tailored for long-term series forecasting, demonstrating state-of-the-art performance while maintaining a streamlined architecture through innovative segmentation and parallel decoding techniques. Nevertheless, SegRNN has several limitations: its fixed segmentation disrupts data continuity and fails to effectively leverage information across different segments, the segmentation strategy employed by SegRNN does not fundamentally address the issue of information loss within the recurrent structure. To address these issues, we propose the MSegRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase. Additionally, we incorporate residual structures in the encoding layer to mitigate information loss within the recurrent structure. To extract information more effectively, we further integrate the Mamba architecture to enhance time series information extraction. Experiments on several real-world long time series forecasting datasets demonstrate that our model surpasses the performance of current state-of-the-art models.
@@ 1 Introduction
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy. Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
@@ 3.1 Overview
In this paper, we propose the MSegRNN method to overcome these limitations through three strategies, as illustrated in Figure 1. We first introduce implicit segmentation, which allows for dense information exchange during the segmentation stage. Subsequently, we incorporate a residual structure to enable partial information bypassing the recurrent structure, thereby reducing information loss in recurrent structure. We also integrate the Mamba architecture to selectively process time series information, further enhancing the model’s capability in handling temporal data.
Figure 1: Overall structure of MSegRNN model
@@ 4 Experiments
Our experiments are structured as follows. We first introduce the datasets used for long-term time series analysis. Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison. Then we presents and analyzes the performance metrics. We further conducted ablation studies to investigate the effectiveness of the model. Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/MSegRNN.git.
@@ 4.3 Main Result
MSegRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
@@
Figure 3: Comparsion of SegRNN and MSegRNN with vary look-back window length
@@ 4.5 Model efficiency
Figure 4: Comparsion of SegRNN and MSegRNN with time and space efficiency.
@@ 5 Conclusion
In this work, we present the MSegRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework. These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model. It is noteworthy that our proposed implicit segmentation method is not limited to a specific long-term time series forecasting method.
We further conducted experiments on six real-world benchmark datasets for long time series. The results demonstrate that our model yields significant imporvements in some real-world scenarios. To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of MSegRNN against SegRNN on datasets of varying scales. The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the MSegRNN method excels with corresponding empirical research.
@@ Appendix C Detail Configuration
Table 4: Detail configuration of MSegRNN
@@ Appendix D Comparison with extra methods
Datasets Models MSegRNN TimeMachine Bi-Mamba4TS DTMamba
@@ Appendix E Prediction Result
Figure 5: Prediction result of different method. The four methods are Dlinear, MSegRNN, SegRNN, and PatchTST in sequence. Intuitively, our method also has the best fitting performance, which is consistent with the fact that MSegRNN has the minimum loss among the models.

### v4

# ISMRNN: An Implicitly Segmented RNN Method with Mamba for Long-Term Time Series Forecasting

@@ Abstract
Long time series forecasting aims to utilize historical information to forecast future states over extended horizons. Traditional RNN-based series forecasting methods struggle to effectively address long-term dependencies and gradient issues in long time series problems. Recently, SegRNN has emerged as a leading RNN-based model tailored for long-term series forecasting, demonstrating state-of-the-art performance while maintaining a streamlined architecture through innovative segmentation and parallel decoding techniques. Nevertheless, SegRNN has several limitations: its fixed segmentation disrupts data continuity and fails to effectively leverage information across different segments, the segmentation strategy employed by SegRNN does not fundamentally address the issue of information loss within the recurrent structure. To address these issues, we propose the ISMRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase. Additionally, we incorporate residual structures in the encoding layer to mitigate information loss within the recurrent structure. To extract information more effectively, we further integrate the Mamba architecture to enhance time series information extraction. Experiments on several real-world long time series forecasting datasets demonstrate that our model surpasses the performance of current state-of-the-art models.
@@ 1 Introduction
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy. Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed ISMRNN, the main contributions of ISMRNN can be summarized as follows:
@@ 3.1 Overview
In this paper, we propose the ISMRNN method to overcome these limitations through three strategies, as illustrated in Figure 1. We first introduce implicit segmentation, which allows for dense information exchange during the segmentation stage. Subsequently, we incorporate a residual structure to enable partial information bypassing the recurrent structure, thereby reducing information loss in recurrent structure. We also integrate the Mamba architecture to selectively process time series information, further enhancing the model’s capability in handling temporal data.
Figure 1: Overall structure of ISMRNN model
@@ 4 Experiments
Our experiments are structured as follows. We first introduce the datasets used for long-term time series analysis. Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison. Then we presents and analyzes the performance metrics. We further conducted ablation studies to investigate the effectiveness of the model. Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/ISMRNN.git.
@@ 4.3 Main Result
ISMRNN SegRNN iTransformer Crossformer PatchTST Dlinear TimesNet
@@
Figure 3: Comparsion of SegRNN and ISMRNN with vary look-back window length
@@ 4.5 Model efficiency
Figure 4: Comparsion of SegRNN and ISMRNN with time and space efficiency.
@@ 5 Conclusion
In this work, we present the ISMRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework. These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model. It is noteworthy that our proposed implicit segmentation method is not limited to a specific long-term time series forecasting method.
We further conducted experiments on six real-world benchmark datasets for long time series. The results demonstrate that our model yields significant imporvements in some real-world scenarios. To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of ISMRNN against SegRNN on datasets of varying scales. The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the ISMRNN method excels with corresponding empirical research.
@@ Appendix C Detail Configuration
Table 4: Detail configuration of ISMRNN
@@ Appendix D Comparison with extra methods
Datasets Models ISMRNN TimeMachine Bi-Mamba4TS DTMamba
@@ Appendix E Prediction Result
Figure 5: Prediction result of different method. The four methods are Dlinear, ISMRNN, SegRNN, and PatchTST in sequence. Intuitively, our method also has the best fitting performance, which is consistent with the fact that ISMRNN has the minimum loss among the models.